OpenStudy (anonymous):

Evaluate 2x^2-3x +1 if x = 3.

OpenStudy (anonymous):

10

OpenStudy (anonymous):

10

OpenStudy (anonymous):

\[2(3)^2 - 3(3) + 1 = 18 - 9 + 1 = 9 + 1 = 10\]